# and Quantization in Prefix Coding

### v6

@@ Abstract
We develop a statistical mechanics framework for prefix coding based on variational principles, renormalization, and quantization. A Lagrangian formulation of the Minimal Entropy principle under the Kraft–McMillan constraint yields a Gibbs-type implied distribution and completeness of the optimal code. A renormalization operator acting on codeword distribution laws produces a coarse-graining flow whose fixed points have iterated-log structure; discrete quantizations of these fixed points include Elias’ $\omega$ code as a special case.
@@ 1 Introduction
Universal integer codes are a classical topic at the interface of information theory and the theory of algorithms. Elias’ family of $\gamma$–, $\delta$–, and $\omega$–codes gives practical schemes whose expected lengths track entropy without prior knowledge of the source law – a property known as universality [1].
@@
Let $s=\text{bin}(x)$ be the binary string representation of a positive decimal number $x$, and let $x=\text{dec}(s)$ be the decimal value corresponding to a binary string $s$, possibly with leading zeros. Then Elias’ $\omega$ is defined by the following two algorithms [2].
@@
Being prefix-free turns out to be equivalent to the classical Kraft–McMillan [3, 4] conditions on the code word lengths. This condition turns out to be equivalent to a specific bound on the statistical sum with respect to a Gibbs measure.
@@
    a Lagrangian formulation of optimal prefix coding under the Kraft–McMillan conditions [3, 4, 5], and
@@
The Lagrangian approach has been used before by Kafri to analyze Benford’s law [6], and by Visser [7] to derive Zipf’s law from the Maximum Entropy Principle [8, 9].
@@
The encoding problem becomes a transparent toy model where scale-invariance is enforced by information-theoretical constraints (being uniquely decodable) rather than dynamics, linking two methodological pillars of statistical physics: variational principles and renormalization flows [10, 11, 12].
@@ 2.1 Lagrangian Optimization for Codelength
Some interesting and practically common statistical laws appear to follow as optimal solutions of Lagrangians combining together the Minimal Entropy Principle with various combinatorial or physical constraints. One notable example is Visser’s derivation of Zipf’s law in [7]. Another one is Kafri’s work of Benford’s law [6] (see also [13]).
@@
Let us recast the Kraft–McMillan inequality [3, 4] for prefix codes in the continuous form:
$\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\leq 1,$ (6)
@@
$\mathcal{L}(\ell)=\int_{\mathcal{C}}\ell(x)p(dx)-\lambda\left(1-\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\right)\to\min_{\ell},$ (8)
@@
The whole Lagrangian will then minimize the average codelength under the prefix-free constraint. This entropy–optimality interpretation is classical [8, 9, 5].
@@
$r(x)=\frac{dp}{d\nu}(x)=\begin{cases}P(w),&x=w\in W,\\[4.0pt] \rho(x),&x\in\mathrm{supp}\,\rho.\end{cases}$ (9)
@@
$\frac{\delta\mathcal{L}}{\delta\ell}=r(x)-\lambda\exp(-\ell(x))=0\;(\nu\text{-a.e.}),$ (10)
@@
$1=p(\mathcal{C})=\int_{\mathcal{C}}r(x)\nu(dx)=\lambda\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)=\lambda.$ (11)
Indeed, the complementary slackness condition implies that the code so derived has to saturate Kraft’s inequality, and thus be complete (see [14] for fundamentals of convex analysis).
@@
$\ell(x)=\begin{cases}-\log P(w),&x=w\in W,\\ -\log\rho(x),&x\in\mathrm{supp}\,\rho.\end{cases}$ (12)
@@ 2.2 Renormalizing the Codeword Distribution
Let us define the following renormalization operator $\mathcal{R}$ for probabilities that is consistent with exponentiation. Namely, we want to be able to renormalize for $\exp(x)$ if we know the probabilities for $x$. This, on the surface, does not imply the universality property shown in [2], but rather guarantees that the conditional distribution is the same over all orders of magnitude. However, we shall see that universality will readily follow from this simple distributional condition.
The process used below is in essence the one of coarse‐graining by replacing $x$ with $\log x$ followed by renormalizing the resulting probability (or codelength). However, it appears to be technically much simpler than the classical renormalization techniques in physics [10, 11, 12] due to its code-theoretical nature.
@@
$\ell(x)=\log x+o(\log x),\text{ as }x\to\infty,$ (16)
and, moreover, we have that for any $\varepsilon>0$ and all $x\in\mathcal{C}$ large enough $o(\log x)\leq\varepsilon\log x$ and thus
$\int_{\mathcal{C}}o(\log x)p(dx)\leq O(1)+\varepsilon\int_{\mathcal{C}}\log(x)p(dx)\leq O(1)+\varepsilon H(p)=O(1),$ (17)
@@
This in its turn yields
$\mathbb{E}[\ell]=\int_{\mathcal{C}}\ell(x)\,p(dx)=\int_{\mathcal{C}}\log(x)\,p(dx)+O(1)\leq H(p)+O(1),$ (18)
where the last step is Wyner’s inequality [15]. Here,
$H(p)=-\int_{\mathcal{C}}\log p(x)\,p(dx)$ (19)
is the Shannon entropy [16] of $\mathcal{C}$. On the other hand, Shannon’s classical lower bound $\mathbb{E}[\ell]\geq H(p)$ also holds (cf. Section 4.4).
@@ 2.3 Renormalization Flow Universality
$\mathcal{T}[\ell](x)\;=\;\ell\bigl(\log x\bigr)\;+\;\log x.$ (21)
@@
$\ell_{*}(x)=\log x\;+\;\log(\log x)\;+\;\log\bigl(\log(\log x)\bigr)\;+\;\cdots,$ (24)
@@
If we rewrite Elias’ $\omega$ codelength using natural logarithms instead of base-$2$ logarithms111i.e. we measure codelengths in nats rather than in bits., we get
$\ell_{\omega}(x)=\ell_{*}(x)+O(\log^{*}x)=\ell_{*}(x)+o(\log(x)^{\varepsilon}),$ (25)
@@
$\mathcal{T}:\ell\mapsto\ell\circ\log+\log$ (28)
@@ 2.4 Passing to the Discrete Encoding
Naturally, if we now recast our observation into the discrete setting, Elias’ $\omega$ encoding [2] will provide a practical realization with very close properties.
@@
$\ell_{\omega}(n)=\ell_{\omega}(\beta(n)-1)+\beta(n).$ (31)
@@
$I_{k}=\{n\in\mathbb{N}\,|\,\beta(n)=k\}=\{n\in\mathbb{N}\,|\,2^{k-1}\leq n\leq 2^{k}-1\}.$ (32)
@@
$S_{k}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(n)}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(k-1)-k}=2^{-\ell_{\omega}(k-1)-1}.$ (33)
@@
$S=\sum^{\infty}_{k=1}S_{k}=S_{1}+\sum^{\infty}_{k=2}S_{k}=\frac{1}{2}+\sum^{\infty}_{k=2}2^{-\ell_{\omega}(k-1)-1}=\frac{1}{2}+\frac{1}{2}\sum^{\infty}_{k=1}2^{-\ell_{\omega}(k)}=\frac{1+S}{2}.$ (34)
This readily implies that $S=1$. The real rate of convergence is very slow though: in our computations we have reached (see [17])
$\sum^{2^{2^{24}}}_{n=1}2^{-\ell_{\omega}(n)}\approx 0.9697265625\ldots$ (35)
@@ 4.1 Mixed Source Distributions
Let $p(x)$ be a probability measure on the set of codewords $\mathcal{C}=[1,+\infty)$ of the form
@@
We say that $\mathcal{C}$ satisfies the Kraft–McMillan condition if
$\int_{\mathcal{C}}\exp(-\ell(x))dx\leq 1,$
where $\ell(x)$ is the codelength of $x\in\mathcal{C}$.
@@ 4.2 Continuous Code Quantization
Then $W_{0}$ is of atoms to encode explicitly, and $\Pi$ is a quantization of the continuous support.
@@ 4.3 Prefix-free Quantized Codes
4.4 Shannon’s Entropy Lower Bound
Let $\mathcal{C}\subset[1,+\infty)$ be a code as defined throughout the manuscript, and let its coarse distribution satisfy
$p(dx)=\sum_{w\in W}\delta_{w}(dx)+\rho(x)dx,$
where $W\cap\mathrm{supp}\,\rho=\emptyset$.
A length function $\ell:\mathcal{C}\to[0,\infty]$ is admissible if
$\int e^{-\ell(x)}\,\nu(dx)\;\leq\;1.$
Then the entropy is
$H(p)=-\sum_{w\in W}P(w)\log P(w)\;-\;\int_{\mathcal{C}}\rho(x)\log\rho(x)\,dx.$
We assume that $\mathcal{C}$ satisfies the Kraft condition:
$K=\int e^{-\ell(x)}\,\nu(dx)\leq 1,$
and call
$q(x)=\frac{e^{-\ell(x)}}{K}$
the Kraft distribution.
The latter is equivalent to
$\ell(x)=-\ln q(x)-\ln K,$
and thus we obtain
$\mathbb{E}_{p}[\ell]$ $=\int_{\mathcal{C}}\ell(x)\,p(dx)=-\int_{\mathcal{C}}\ln q(x)\,p(dx)-\ln Z$
$=H(p)+D_{\mathrm{KL}}(p\|q)-\ln Z\;\geq\;H(p),$
since $D_{\mathrm{KL}}(p\|q)\geq 0$ (by Gibbs’ inequality) and $\ln Z\geq 0$ (by assumption).
Thus, we have an analogue of Shannon’s lower entropy bound
$\mathbb{E}_{p}[\ell]\ \geq\ H(p),$
with equality if and only if $K=1$ and $e^{-\ell}=\frac{dp}{d\nu}$ ($\nu$-a.e.).
@@ 4.5 Heisenberg-style and Boltzmann-style relations
is the Kullback–Leibler (KL) divergence of two distributions [5].
@@
$\bar{\ell}=\big(H(p)-\sum_{j=1}^{m}\pi_{A_{j}}\ln v_{j}\big)\;+\;D_{\mathrm{KL}}(\rho\|\bar{\rho})\;+\;D_{\mathrm{KL}}(\pi\|q)\;+\;\ln\frac{1}{K},$
@@
    The leading term $H(p)-\sum_{j}\pi_{A_{j}}\ln v_{j}$ is the “resolution-adjusted entropy.”
@@
    The nonnegative correction terms reflect: within-cell mismatch $D_{\mathrm{KL}}(\rho\|\bar{\rho})$, Huffman tree mismatch $D_{\mathrm{KL}}(\pi\|q)$, and Kraft slack $\ln(1/K)$.
@@
For a uniform distribution on a region $U\subset\mathbb{R}$ of measure $V$, the mixed entropy is
$H(p)=h(\rho)=\ln V,$
exactly mirroring Boltzmann’s relation $S=k_{B}\ln\Omega$, with $\Omega$ being the number (or volume) of accessible microstates. Then $e^{H(p)}$ is the perplexity in information theory and the multiplicity of microstates in statistical physics.
@@
$V_{Q}=\exp\!\Big(\sum_{j}\pi_{A_{j}}\ln v_{j}\Big).$
This plays the role of a coarse-graining scale: the effective cell size in which the source is observed. In physics, such a scale is set by Planck’s constant $h^{d}$ when passing from classical to quantum phase space.

### v7

@@ Abstract
We develop a statistical mechanics framework for prefix coding based on variational principles, renormalization, and quantization. A Lagrangian formulation of entropy-optimal encoding under the Kraft–McMillan constraint yields a Gibbs-type implied distribution and completeness of the optimal code. A renormalization operator acting on codeword distribution laws produces a coarse-graining flow whose fixed points have iterated-log structure; discrete quantizations of these fixed points include Elias’ $\omega$ code as a special case.
@@ 1 Introduction
Universal integer codes are a classical topic at the interface of information theory and the theory of algorithms. Elias’ family of $\gamma$–, $\delta$–, and $\omega$–codes gives practical schemes whose expected lengths track entropy without prior knowledge of the source law – a property known as universality [undef].
@@
Let $s=\text{bin}(x)$ be the binary string representation of a positive decimal number $x$, and let $x=\text{dec}(s)$ be the decimal value corresponding to a binary string $s$, possibly with leading zeros. Then Elias’ $\omega$ is defined by the following two algorithms [undefa].
@@
Being prefix-free turns out to be equivalent to the classical Kraft–McMillan [undefb, undefc] conditions on the code word lengths. This condition turns out to be equivalent to a specific bound on the statistical sum with respect to a Gibbs measure.
@@
    a Lagrangian formulation of optimal prefix coding under the Kraft–McMillan conditions [undefb, undefc, undefd], and
@@
The Lagrangian approach has been used before by Kafri to analyze Benford’s law [undefe], and by Visser [undeff] to derive Zipf’s law from the Maximum Entropy Principle [undefg, undefh].
@@
The encoding problem becomes a transparent toy model where scale-invariance is enforced by information-theoretical constraints (being uniquely decodable) rather than dynamics, linking two methodological pillars of statistical physics: variational principles and renormalization flows [undefi, undefj, undefk].
@@ 2.1 Lagrangian Optimization for Codelength
Some interesting and practically common statistical laws appear to follow as optimal solutions of Lagrangians combining together the entropy-optimal encoding condition with various combinatorial or physical constraints. One notable example is Visser’s derivation of Zipf’s law in [undeff]. Another one is Kafri’s work of Benford’s law [undefe] (see also [undefl]).
@@
Let us recast the Kraft–McMillan inequality [undefb, undefc] for prefix codes in the continuous form:
$\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\leq 1,$
@@
$\mathcal{L}(\ell)=\int_{\mathcal{C}}\ell(x)p(dx)-\lambda\left(1-\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)\right)\to\min_{\ell},$
@@
The whole Lagrangian will then minimize the average codelength under the prefix-free constraint. This entropy–optimality interpretation is classical [undefg, undefh, undefd].
@@
$r(x)=\frac{dp}{d\nu}(x)=\begin{cases}P(w),&x=w\in W,\\[4.0pt] \rho(x),&x\in\mathrm{supp}\,\rho.\end{cases}$
@@
$\frac{\delta\mathcal{L}}{\delta\ell}=r(x)-\lambda\exp(-\ell(x))=0\;(\nu\text{-a.e.}),$
@@
$1=p(\mathcal{C})=\int_{\mathcal{C}}r(x)\nu(dx)=\lambda\int_{\mathcal{C}}\exp(-\ell(x))\nu(dx)=\lambda.$
Indeed, the complementary slackness condition implies that the code so derived has to saturate Kraft’s inequality, and thus be complete (see [undefm] for fundamentals of convex analysis).
@@
$\ell(x)=\begin{cases}-\log P(w),&x=w\in W,\\ -\log\rho(x),&x\in\mathrm{supp}\,\rho.\end{cases}$
@@ 2.2 Renormalizing the Codeword Distribution
Let us define the following renormalization operator $\mathcal{R}$ for probabilities that is consistent with exponentiation. Namely, we want to be able to renormalize for $\exp(x)$ if we know the probabilities for $x$. This, on the surface, does not imply the universality property shown in [undefa], but rather guarantees that the conditional distribution is the same over all orders of magnitude. However, we shall see that universality will readily follow from this simple distributional condition.
The process used below is in essence the one of coarse‐graining by replacing $x$ with $\log x$ followed by renormalizing the resulting probability (or codelength). However, it appears to be technically much simpler than the classical renormalization techniques in physics [undefi, undefj, undefk] due to its code-theoretical nature.
@@
$\ell(x)=\log x+o(\log x),\text{ as }x\to\infty.$ (4)
Recall that the codeword probability $p(x)$ is implied, i.e. $r(x)=\frac{dp}{d\nu}(x)=e^{-\ell(x)}$ $(\nu$-a.e.). From equations (3) – (4) we have that $\ell(x)\geq\log x$, as the additional polylog terms in the expression for $\ell(x)$ only arise while they are non-negative. Hence
$r(x)=e^{-\ell(x)}\leq e^{-\log x}=\frac{1}{x},$
so $x\leq 1/r(x)$ and therefore $\log x\leq-\log r(x)$ pointwise.
In turn, this yields
$\int_{\mathcal{C}}\log x\,p(dx)\leq H(p),$
where
$H(p)=-\int_{\mathcal{C}}\log r(x)\,p(dx)$
is the Shannon entropy [undefn] of $\mathcal{C}$.
This implies that for any $\varepsilon>0$ and all $x\in\mathcal{C}$ large enough $o(\log x)\leq\varepsilon\log x$ and thus
$\int_{\mathcal{C}}o(\log x)p(dx)\leq O(1)+\varepsilon\int_{\mathcal{C}}\log x\,p(dx)\leq O(1)+\varepsilon H(p)=O(1),$
@@
Finally,
$\mathbb{E}[\ell]=\int_{\mathcal{C}}\ell(x)\,p(dx)=\int_{\mathcal{C}}\log x\,p(dx)+O(1)\leq H(p)+O(1).$
On the other hand, Shannon’s classical lower bound $\mathbb{E}[\ell]\geq H(p)$ also holds (cf. Section 4.2).
@@ 2.3 Renormalization Flow Universality
$\mathcal{T}[\ell](x)\;=\;\ell\bigl(\log x\bigr)\;+\;\log x.$
@@
$\ell_{*}(x)=\log x\;+\;\log(\log x)\;+\;\log\bigl(\log(\log x)\bigr)\;+\;\cdots,$
@@
If we rewrite Elias’ $\omega$ codelength using natural logarithms instead of base-$2$ logarithms333i.e. we measure codelengths in nats rather than in bits., we get
$\ell_{\omega}(x)=\ell_{*}(x)+O(\log^{*}x)=\ell_{*}(x)+o(\log(x)^{\varepsilon}),$
@@
$\mathcal{T}:\ell\mapsto\ell\circ\log+\log$
@@ 2.4 Passing to the Discrete Encoding
Naturally, if we now recast our observation into the discrete setting, Elias’ $\omega$ encoding [undefa] will provide a practical realization with very close properties.
@@
$\ell_{\omega}(n)=\ell_{\omega}(\beta(n)-1)+\beta(n).$
@@
$I_{k}=\{n\in\mathbb{N}\,|\,\beta(n)=k\}=\{n\in\mathbb{N}\,|\,2^{k-1}\leq n\leq 2^{k}-1\}.$
@@
$S_{k}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(n)}=\sum_{n\in I_{k}}2^{-\ell_{\omega}(k-1)-k}=2^{-\ell_{\omega}(k-1)-1}.$
@@
$S=\sum^{\infty}_{k=1}S_{k}=S_{1}+\sum^{\infty}_{k=2}S_{k}=\frac{1}{2}+\sum^{\infty}_{k=2}2^{-\ell_{\omega}(k-1)-1}=\frac{1}{2}+\frac{1}{2}\sum^{\infty}_{k=1}2^{-\ell_{\omega}(k)}=\frac{1+S}{2}.$
This readily implies that $S=1$. The real rate of convergence is very slow though: in our computations we have reached (see [undefo])
$\sum^{2^{2^{24}}}_{n=1}2^{-\ell_{\omega}(n)}\approx 0.9697265625\ldots$
@@ 4.1 Mixed Source Distributions
Let $p(x)$ be a probability measure on the set of codewords $\mathcal{C}\subseteq[1,+\infty)$ of the form
@@
4.2 Shannon’s Entropy Lower Bound
The Shannon entropy of $\mathcal{C}$ is
$H(p)=-\int_{\mathcal{C}}\log\left(\frac{dp}{d\nu}(x)\right)p(dx)=-\sum_{w\in W}P(w)\log P(w)\;-\;\int_{\mathcal{C}}\rho(x)\log\rho(x)\,dx,$
where
$\nu(dx)=\sum_{w\in W}\delta_{w}(dx)+dx$
is the Kraft measure.
We assume that $\mathcal{C}$ satisfies the Kraft–McMillan condition:
$K=\int e^{-\ell(x)}\,\nu(dx)\leq 1,$
and call
$q(x)=\frac{e^{-\ell(x)}}{K}$
the Kraft distribution.
The latter is equivalent to
$\ell(x)=-\log q(x)-\log K,$
and thus we obtain
$\mathbb{E}_{p}[\ell]$ $=\int_{\mathcal{C}}\ell(x)\,p(dx)=-\int_{\mathcal{C}}\log q(x)\,p(dx)-\log K$
$=H(p)+D_{\mathrm{KL}}(p\|q)-\log K\;\geq\;H(p),$
since $D_{\mathrm{KL}}(p\|q)\geq 0$ (by Gibbs’ inequality) and $\log K\geq 0$ (by assumption).
Thus, we have an analogue of Shannon’s lower entropy bound
$\mathbb{E}_{p}[\ell]\ \geq\ H(p),$
with equality if and only if $K=1$ and $e^{-\ell}=\frac{dp}{d\nu}$ ($\nu$-a.e.).
@@ 4.3 Continuous Code Quantization
Then $W_{0}$ is a set of atoms to encode explicitly, and $\Pi$ is a quantization of the continuous support.
@@ 4.5 Heisenberg-style and Boltzmann-style relations
is the Kullback–Leibler (KL) divergence of two distributions [undefd].
@@
$\bar{\ell}=\big(H(p)-\sum_{j=1}^{m}\pi_{A_{j}}\log v_{j}\big)\;+\;D_{\mathrm{KL}}(\rho\|\bar{\rho})\;+\;D_{\mathrm{KL}}(\pi\|q)\;+\;\log\frac{1}{K},$
@@
    The leading term $H(p)-\sum_{j}\pi_{A_{j}}\log v_{j}$ is the “resolution-adjusted entropy.”
@@
    The nonnegative correction terms reflect: within-cell mismatch $D_{\mathrm{KL}}(\rho\|\bar{\rho})$, Huffman tree mismatch $D_{\mathrm{KL}}(\pi\|q)$, and Kraft slack $\log(1/K)$.
@@
$V_{Q}=\exp\!\Big(\sum_{j}\pi_{A_{j}}\log v_{j}\Big).$
This plays the role of a coarse-graining scale: the effective cell size in which the source is observed. In physics, such a scale is set by Planck’s constant when passing from classical to quantum phase space.